Find the general solution of following

1) y''-2y'+y=exsin-1(x)



1
Expert's answer
2021-08-10T13:53:13-0400

"\displaystyle
y'' - 2y' + y = e^x\arcsin{x}\\
\textsf{The auxiliary equation is}\\
m^2 - 2m + 1 = 0\\
(m - 1)^2 = 0\\
m = 1 \,\, \textsf{twice}\\
y = e^x(C_1 + xC_2)\\
\textsf{Wronskian}(e^x, xe^x) = e^{2x}\\
\textsf{Let} \,\, C_1 \,\,\textsf{and}\,\, C_2 \,\, \textsf{be}\,\, \textsf{functions}\\
K_1 \,\,\textsf{and}\,\, K_2.\\
K_1 = - \int \frac{xe^x \times e^{x}\arcsin{x}}{e^{2x}}\mathrm{d}x = -\int x\arcsin{x}\mathrm{d}x \\

\textsf{and} \,\, K_2 = \int \frac{e^x \times e^{x}\arcsin{x}}{e^{2x}}\mathrm{d}x = \int \arcsin{x}\mathrm{d}x \\

\textsf{For}\,\, K_1, \,\, \textsf{Let}\,\, x = \sin{u}\\
\implies \mathrm{d}(\sin{u}) = \cos{u}\,\,\mathrm{d}u\\
\begin{aligned}
K_1 &= -\int x\arcsin{x}\mathrm{d}x = -\int u\sin{u}\cos{u}\mathrm{d}u
\\&= -\frac{1}{2}\int u\sin{2u}\mathrm{d}u
\\&= \frac{1}{2}\int u\mathrm{d}\left(\frac{\cos{2u}}{2}\right)
\\&= \frac{u\cos{2u}}{4} - \frac{1}{4}\int \cos{2u}\mathrm{d}u
\\&= \frac{u\cos{2u}}{4} - \frac{\sin{2u}}{8} + C
\\&= \frac{\arcsin{x}\cos{\left(2\arcsin{x}\right)}}{4} - \frac{\sin\left(2\arcsin{x}\right)}{8} + C_1
\end{aligned}\\

\textsf{For}\,\, K_2 \,\, \textsf{Let}\,\, x = \sin{u}\\
\implies \mathrm{d}(\sin{u}) = \cos{u}\,\,\mathrm{d}u\\
\begin{aligned}
K_2 &= \int \arcsin{x}\mathrm{d}x = \int u\cos{u}\mathrm{d}u
\\&= u\sin{u} - \int \sin{u}\mathrm{d}u
\\&= u\sin{u} + \cos{u} + C
\\&= x\arcsin{x} + \cos\left(\arcsin{x}\right) + C_2
\end{aligned}\\

\textsf{But}\,\,\, y = e^x K_1 + xe^xK_2\\

\begin{aligned}
\implies y &= e^x\left(\frac{\arcsin{x}\cos{\left(2\arcsin{x}\right)}}{4} -\frac{\sin\left(2\arcsin{x}\right)}{8} + C_1\right)
\\&\hspace{2cm}+ xe^x\left(x\arcsin{x} + \cos\left(\arcsin{x}\right) + C_2\right)
\\&= C_1e^x + C_2xe^x + \frac{e^x\arcsin{x}\cos{\left(2\arcsin{x}\right)}}{4} - \frac{e^x\sin\left(2\arcsin{x}\right)}{8}
\\&\hspace{2cm}+ x^2e^x\arcsin{x} + xe^x\cos\left(\arcsin{x}\right)
\end{aligned}\\

\textsf{Note that the following holds}\\
\cos(2\arcsin(x)) = 1 - 2x^2\\
\sin(2\arcsin(x)) = 2x\sqrt{1 - x^2}\\
\cos(\arcsin(x)) = \sqrt{1 - x^2}\\

\textsf{Substituting the above in}\,\,\, y\\
\begin{aligned}
y &= C_1e^x + C_2xe^x + \frac{e^x\arcsin{x}(1 - 2x^2)}{4} - \frac{2xe^x\sqrt{1 - x^2}}{8}
\\&\hspace{2cm}+ x^2e^x\arcsin{x} + xe^x\sqrt{1 - x^2}
\\&= C_1e^x + C_2xe^x + \frac{e^x\arcsin{x}}{4} - \frac{x^2e^x\arcsin{x}}{2} - \frac{xe^x\sqrt{1 - x^2}}{4}
\\&\hspace{2cm}+ x^2e^x\arcsin{x} + xe^x\sqrt{1 - x^2}
\\&= C_1e^x + C_2xe^x + \frac{e^x\arcsin{x}}{4} + \frac{x^2e^x\arcsin{x}}{2} + \frac{3xe^x\sqrt{1 - x^2}}{4}
\end{aligned}\\

\implies y = C_1e^x + C_2xe^x + \frac{3xe^x\sqrt{1 - x^2}}{4}+ \frac{x^2e^x\arcsin{x}}{2} + \frac{e^x\arcsin{x}}{4}"